A coil of inductance 5.0 mH and negligible resistance is connected to the oscillator of the previous problem. Find the peak currents in the circuit for ω = 100 s^-1, 500 s^-1, 1000 s^-1.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. Inductance = 5.0 mH = 0.005 H a) ω = 100 s^–1 X base L = ωL = 100 * 5/1000 = 0.5 Ω i = ε base 0/X base L = 10/0.5 = 20 A b) ω = 500 s^–1 X base L = ωL = 500 * 5/1000 = 2.5 Ω i = ε base 0/X base L = 10/2.5 = 4 A c) ) ω = 1000 s^-1 X base L = ωL = 1000 * 5/1000 = 5 Ω i = ε base 0/X base L = 10/5 = 2 A